OpenStudy (anonymous):

20 + 25 + 50(x - 2) = 145

OpenStudy (anonymous):

wouldn't i add 20+25+50 and get 95 then distribute it to the x and 2?

OpenStudy (anonymous):

45+ 50X-100=145 subtract 45 from both sides and add 100 to both sides 50x=200 divide both sides by 50 x=4

OpenStudy (helpsos):

\[20+ 25+ 50(x-20) = 145\] \[20 + 25 + 50x - 100 = 145\] \[45 + 50x -100 = 145\] \[-55 + 50x = 145\] \[50x = 200\] \[x = 4\]

OpenStudy (anonymous):

No, because the 50 is being multiplied by the (x-2) so you need to distribute the 50 first, and then you can add. Remember order of operations (PEMDAS) Parenthesis Exponents Multiply Divide Add Subtract. So you would multiply before you add

OpenStudy (anonymous):

ohhhhh okay i didn't think about that

OpenStudy (helpsos):

sorry I just wrote the equation but I hoped it helped some :)

OpenStudy (anonymous):

yeah it helped a lot

OpenStudy (anonymous):

I'm super tired right now

OpenStudy (anonymous):

Ya, no worries, I used to do that all of the time. A good way to remember this is Please Excuse My Dear Aunt Sally. (PEMDAS)

OpenStudy (anonymous):

or if you can find another one you like :)

OpenStudy (anonymous):

thank you for your help :)

OpenStudy (anonymous):

You're welcome...good luck!